OpenStudy (anonymous):

peter bought 6 mangoes at $x each. He also bought a papaya for $5. a) express the total cost of the fruits in terms of x. b) if x=2, find the total cost of the fruits.

OpenStudy (anonymous):

a) cost = 6x + 5 b) cost = 6(2) + 5 = 12 + 5 = $17